You are given a system where accessing memory takes 500 cycles per memory access. A hit in the
cache takes 10 cycles and the latency of a cache miss (including the time to search in the cache and
getting the data from memory) is 500 cycles.


• (2 points) You run the program by completely removing the caches. How many cycles does the
program take?


• (2 points) You run the program with caches and 90% of the accesses are hits in the cache and
the rest are misses. How many cycles will the program take with caches?


• (2 points) Assume that your program takes 10 hours to run on a machine without caches. How
long would it take on a machine with caches?
Answer #1

memory access = 500 cycles
cache access = 10 cycles
latency of cache miss = 500 cycles

(a)Without cache memory ,it need to access memory for each instruction fetch.
cycles needed per instruction = 500 cycles

(b)cache hit rate = 90%
cycles needed per instruction = 0.9 * 10 + 0.1* 500 = 9 + 50 = 59 cycles

(c)without cache it takes 10 hours
with cache it will take = (10/500)*59 = 1.18 hours